# A “Staircase” formula for the Chern-Schwartz-MacPherson cycle of a matroid

### v2

@@ 1 Introduction
De Concini and Procesi defined the notion of a wonderful model of a hyperplane arrangement with respect to a building set in [DCP96]. In this paper, we will use the term wonderful model of $\mathcal{A}$ to refer to the wonderful model with respect to the maximal building set. Later, Feichtner and Yuzhvinsky [FY04] define the Chow ring of a matroid following the ideas of [DCP96] and the correspondence between the Chow ring of a nonsingular proper toric variety $X$ and Minkowski weights on its associated fan $\Sigma$ proven by Fulton and Strumfels in [FS97]. In [AHK18], Adiprasito, Huh and Katz bring these concepts together and prove that $A^{*}(M)$ and $MW_{*}(B(M))$ are dual to each other, in analogy to Poincaré duality for the toric case and the realizable case. This gave the necessary setting for López de Medrano, Rincón, and Shaw to define an analogous CSM cycle for matroids [LdMRS19].
@@
where $M$ is the matroid associated to $\mathcal{A}$.
@@
indexed over all size $k$ multisubset chains =$\mathcal{F}=\{F_{r_{1}}\subseteq\dots\subseteq F_{r_{k}}\}$ of flats of $M$, with coefficients $C_{\mathcal{F}}$ on each monomial
@@
the staircase class of a matroid, which generalizes the formula provided in Section 3 to the case where $M$ is an arbitrary matroid and prove in Theorem 4.1 that this class satisfies a contraction-deletion formula similar to that of $\operatorname{csm}(M)$. The name comes from the visual structure of the class when written out explicitly; in the case where $M=U_{3,3}$ with underlying set $E=\{0,1,2,3\}$,
@@
The proof of this statement uses the contraction-deletion formula, the interpretation of the Chow ring in terms of piecewise polynomials [Bri96] and induction on the number of elements of $M$, which completes the proof of Conjecture 5.2.4 in [Man22].
Other work: In [Rau23], Rau provides a formula for the CSM cycle of a matroid $M$ in terms of the tropical polynomials cutting out the Bergman fan $1_{M}$, which is similar to Lemma 3.2 if we interpret the tropical polynomials as the sections whose zero sections are the components of the pullback of the hyperplane arrangement in the realizable case. In [BEST23], Berget, Eur, Spink and Tseng construct the framework of tautological classes of matroids, which generalizes the class of the universal sub and quotient bundle of the Grassmannian to an arbitrary matroid. Through this framework they obtain a formula for the inclusion of the group of Minkowski weights of a matroid $M$ into the Chow ring of the corresponding permutahedral variety and show that the Poincaré dual of $\operatorname{csm}(M)$ in the Chow ring of the permutahedral variety is $\sum_{k=0}^{\operatorname{rk}(M)-1}c_{k}(\mathcal{S_{M}})$. Although our proof of Theorem 4.1 does not use this framework, our result gives the Chern roots of $[\mathcal{S}_{M}]$ as $\{-\sum_{\begin{subarray}{c}F\in\mathcal{L}\\ \operatorname{rk}(F)\geq r\end{subarray}}x_{F}|1\leq r\leq\operatorname{rk}(M)\}$ in $A^{*}(M)$, via appendix III in [BEST23].
@@ 2 Background
Definition 2.3 (Contraction and Deletion).
Given an element $i\in E$ and a matroid $M=(E,\operatorname{rk})$ on $E$, we define the deletion $M\setminus i$ as the matroid $(E\setminus\{i\},\operatorname{rk}|_{2^{E\setminus\{i\}}})$ and the contraction $M/i$ as the matroid $(E\setminus\{i\},\operatorname{rk}^{\prime})$ with $\operatorname{rk}^{\prime}(A)=\operatorname{rk}(A\cup\{i\})-\operatorname{rk}(\{i\})$.
@@
Matroids arising in this way from a hyperplane arrangement in $\mathbb{P}^{d}_{\mathbb{C}}$ (or a vector arrangement in $\mathbb{C}^{d+1}$) are called $\mathbb{C}$-realizable. Note that there exists matroids which are not realizable.
In the realizable case, if $M=(\{0,\dots,k\},\operatorname{rk})$ is represented by the set of vectors $\{v_{0},\dots,v_{k}\}$ in $\mathbb{C}^{n}$, then $M\ i$ is represented by the set of vectors $\{v_{0},\dots,v_{k}\}\setminus\{v_{i}\}$ in $\mathbb{C}^{n}$ and $M/i$ is represented by $\{v_{0}+\braket{v_{i}},\dots,v_{k}+\braket{v_{i}}\}\setminus\{\braket{v_{i}}\}$ in $\mathbb{C}^{n}/\braket{v_{i}}$.
@@
Let $\pi:\mathcal{W}\rightarrow\mathbb{P}^{d}$ denote this blow-up map. $\pi$ is an isomorphism on the complement of the arrangement $C(\mathcal{A})$. On the $\mathcal{W}$ there is a divisor corresponding to each flat of the matroid. All collections of these divisors intersect transversely if their flats are comparable, and are disjoint otherwise. See [DCP96] for details.
@@
for all $i\in E$.
When a matroid is $\mathbb{C}$-realizable, arising from a hyperplane arrangement $\mathcal{A}$, its Chow ring is isomorphic to the Chow ring of the wonderful model $A^{*}(\mathcal{W})$ in the geometric sense, with each $x_{F}$ corresponding to the class of the divisor associated to the proper flat $F$, and $x_{E}=-\pi^{*}(h)$, minus the pullback of the hyperplane class in $\mathbb{P}^{d}$.
@@
Let $M$ be the matroid in Example 2.1 Its Chow ring is presented by the quotient ring
@@
Let $E$ be a finite set and for $S\subset E$ define $e_{S}$ be the element of $\mathbb{R}^{E}$ given by $e_{E}(i)=1$ for all $i\in S$. Given a matroid $M=(E,\operatorname{rk})$, we define $\hat{B}(M)$, the affine Bergman fan of $M$ as the fan in $\mathbb{R}^{E}$ whose cones are of the form
@@
where =$\mathcal{F}=\{\emptyset\subsetneq F_{1}\subsetneq\dots\subsetneq F_{k}\subset E\}$ is a flag of non-empty flats of $M$. We define the image of this fan under the projection $\mathbb{R}^{E}\to\mathbb{R}^{E}/\mathbb{R}e_{E}$ to be $B(M)$, the Bergman fan of $M$. Note that this projection maps $\hat{\sigma}_{\mathcal{F}}$ and $\hat{\sigma}_{\mathcal{F}\cup\{E\}}$ to the same cone, where $\mathcal{F}$ is a flag of proper non-empty flats. We denote the cone corresponding to $\mathcal{F}$ in $B(M)$ by $\sigma_{\mathcal{F}}$.
@@
where $v_{\sigma/\tau}$ is the generator of $\sigma+\braket{\tau}$ in $N_{\mathbb{R}}/\braket{\tau}$ Given a fan $\Sigma$ of pure dimension $n$, we define $k$-dimensional Minkowski weights on $\Sigma$ to be the set of balanced weighted subfans of $\Sigma$.
@@
Given a pure $n$-dimensional, balanced, unimodular fan $\Sigma$ in $N_{\mathbb{R}}$, $A^{k}(\Sigma)\cong MW_{n-k}(\Sigma)$. We denote the isomorphism by _$\_\cap 1_{\Sigma}$, where $1_{\Sigma}$ is the function $\sigma\mapsto 1$ for all $\sigma\in\Sigma(n)$. In the case where $\Sigma=B(M)$ for some matroid $M$, we simply write $1_{M}$ for $1_{B(M)}$. In particular, Remark 2.2 and the statement above imply that for all matroids $M$, $MW_{*}(\hat{B}(M))\cong MW_{*}(B(M))$.
@@ 3 A “Staircase” CSM Class Formula for Realizable Matroids
López de Medrano, Rincón, and Shaw defined an analogous CSM cycle for matroids, and proved that when a matroid is realizable over $\mathbb{C}$, its CSM cycle corresponds to the geometric CSM class $c_{SM}(\mathcal{W}\setminus\mathcal{A})$ of the complement of the arrangement in the wonderful model under the identification of $A^{*}(M)\cong A^{*}(\mathcal{W})$ (Theorem 3.1 in [LdMRS19]).
@@
In this section, we prove such a formula for the CSM class of the complement of the arrangement by geometric means, thus resolving the conjecture for all realizable matroids.
@@
(Lemma 1.3 in [Auf09]) Let $X\subseteq Y$ be a nonsingular varieties, and $\pi:\tilde{Y}\rightarrow Y$ the blow-up of Y along $X$, and $\tilde{X}$ the exceptional divisor. If $X$ is a complete intersection of hypersurfaces $Z_{1},\dots,Z_{d}$ meeting transversally in $Y$, then the Chern class of the tangent bundle of the blow up $\tilde{Y}$ is related to $c(T_{Y})$ by the formula
@@
The wonderful model $\mathcal{W}$ is constructed by a sequence of blow-ups of $\mathbb{P}^{d}$ at the intersection of hyperplanes determined by each flat $F\in\hat{\mathcal{L}}$. If we blow-up according to order by rank, highest to lowest (i.e. by dimension, lowest to highest), the locus of blow-up is intersection $\operatorname{rk}(F)$ many hyperplanes at the intersection determined by $F$ minus the exceptional divisors already blown up in the intersection (corresponding to the flats $F^{\prime}$ which contain $F$. That is, when all blow-ups are complete, each of the $\operatorname{rk}(F)$ hyperplanes in locus of blow-up pulls back to
$\pi^{*}h-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}$
in $A^{*}\mathcal{W}$, where $h$ is the class of the hyperplane in $\mathbb{P}^{d}$. By applying the blow-up formula (Lemma 3.1) according to each flat $F\in\hat{\mathcal{L}}$,
$c(T_{\mathcal{W}})=\prod_{F\in\hat{\mathcal{L}}}\left[\frac{(1+x_{F})(1+(\pi^{*}h-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}})-x_{F})^{\operatorname{rk}(F)}}{(1+(\pi^{*}h-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}))^{\operatorname{rk}(F)}}\right]\pi^{*}c(T_{\mathbb{P}^{d}}).$
@@
$=\frac{1}{\prod_{F\in\hat{\mathcal{L}}}(1+x_{F})}\prod_{F\in\hat{\mathcal{L}}}\left[\frac{(1+x_{F})(1+(\pi^{*}h-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}})-x_{F})^{\operatorname{rk}(F)}}{(1+(\pi^{*}h-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}))^{\operatorname{rk}(F)}}\right]\pi^{*}c(T_{\mathbb{P}^{d}})\cap[\mathcal{W}].$
By cancellation, and $\pi^{*}c(T_{\mathbb{P}^{d}})=(1+\pi^{*}h)^{d+1}=(1+\pi^{*}h)^{\operatorname{rk}(E)}$, we get
$c_{SM}(\mathcal{W}\setminus\mathcal{A})=\dots$ $=\prod_{F\in\hat{\mathcal{L}}}\left[\frac{(1+(\pi^{*}h-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}})-x_{F})^{\operatorname{rk}(F)}}{(1+(\pi^{*}h-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}))^{\operatorname{rk}(F)}}\right]\pi^{*}c(T_{\mathbb{P}^{d}})\cap[\mathcal{W}]$
$=\prod_{F\in\hat{\mathcal{L}}}\left[\frac{(1+(\pi^{*}h-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}})-x_{F})^{\operatorname{rk}(F)}}{(1+(\pi^{*}h-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}))^{\operatorname{rk}(F)}}\right](1+\pi^{*}h)^{\operatorname{rk}(E)}\cap[\mathcal{W}].$
We can substitute $x_{E}:=-\pi^{*}h$ to simplify notation:
$c_{SM}(\mathcal{W}\setminus\mathcal{A})=\dots$ $=\prod_{F\in\hat{\mathcal{L}}}\left[\frac{(1+(-x_{E}-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}})-x_{F})^{\operatorname{rk}(F)}}{(1+(-x_{E}-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}))^{\operatorname{rk}(F)}}\right](1-x_{E})^{\operatorname{rk}(E)}\cap[\mathcal{W}]$
@@
$\mathcal{F}$ must be a chain, since otherwise $\prod_{F\in\mathcal{F}}x_{F}=0$ has empty intersection in $A^{*}\mathcal{W}$. Choose a maximal flag $\{\emptyset=F_{0}\subsetneq F_{1}\subsetneq\dots\subsetneq F_{d}\subsetneq F_{d+1}=E\}$ (where $\operatorname{rk}(F_{i})=i$) containing the chain $\mathcal{F}$ as a multisubset. We may rewrite the monomial as
@@
For all $F$ not included in the flag, we may set $x_{F}=0$. This has no effect on $\prod_{F\in\mathcal{F}}x_{F}$, and only kills some other monomials). If $F$ is not in the flag, the fraction $\left(\frac{1-\sum_{F^{\prime}\supseteq F}x_{F^{\prime}}}{1-\sum_{F^{\prime}\supsetneq F}x_{F^{\prime}}}\right)^{\operatorname{rk}(F)}$ becomes $1$. Otherwise, when $F=F_{r}$ in the flag, it becomes $\left(\frac{1-\sum_{i=r}^{\operatorname{rk}(E)}x_{F_{i}}}{1-\sum_{i=r+1}^{\operatorname{rk}(E)}x_{F_{i}}}\right)^{r}$. The expression then simplifies by cancellation (we indicate the result of killing the non-flag divisors by $\mapsto$):
$\prod_{F\in\mathcal{L}}\left(\frac{1-\sum_{F^{\prime}\supseteq F}x_{F^{\prime}}}{1-\sum_{F^{\prime}\supsetneq F}x_{F^{\prime}}}\right)^{\operatorname{rk}(F)}$ $\mapsto\prod_{r=1}^{\operatorname{rk}(E)}\left(\frac{1-\sum_{i=r}^{\operatorname{rk}(E)}x_{F_{i}}}{1-\sum_{i=r+1}^{\operatorname{rk}(E)}x_{F_{i}}}\right)^{r}$ $=\prod_{r=1}^{\operatorname{rk}(E)}\left(\frac{1-(x_{F_{r}}+x_{F_{r+1}}+\dots+x_{E})}{1-(x_{F_{r+1}}+\dots+x_{E})}\right)^{r}$
$=\prod_{r=1}^{\operatorname{rk}(E)}\left(1-\sum_{i=r}^{\operatorname{rk}(E)}x_{F_{i}}\right)$ $=\prod_{r=1}^{\operatorname{rk}(E)}(1-(x_{F_{r}}+\dots+x_{E})).$
The coefficient of the monomial $x_{F_{r_{1}}}\cdots x_{F_{r_{k}}}$ may be deduced by counting how many times it the appears in the expansion.111alternatively, we can use the Taylor’s Theorem to compute the coefficient by taking the appropriate partial derivatives at zero. The count amounts to choosing from which factor each $x_{F_{r_{i}}}$ comes, then dividing to account for identical permutations. Note that each $-x_{F_{r_{i}}}$ appears as a term only in the first $r_{i}$ factors. There are $r_{1}$ possible choices for $x_{F_{r_{1}}}$, then $r_{2}-1$ many remaining choices for $x_{F_{r_{2}}}$, and so on (i.e. $r_{i}-i+1$ choices for $x_{F_{r_{i}}}$ once the previous $i-1$ variables have been chosen). Therefore, the coefficient on $x_{F_{r_{1}}}\cdots x_{F_{r_{k}}}$ is
@@
The formula conjectured by Fife and Rincón holds for realizable matroids.

### v3

@@ 1 Introduction
De Concini and Procesi defined the notion of a wonderful model of a hyperplane arrangement with respect to a building set in [DCP96]. In this paper, we will use the term wonderful model of $\mathcal{A}$ to refer to the wonderful model with respect to the maximal building set. Later, Feichtner and Yuzhvinsky [FY04] define the Chow ring $A^{*}(M)$ of a matroid $M$ following the ideas of [DCP96] and the correspondence between the Chow ring of a nonsingular proper toric variety $X$ and Minkowski weights $MW_{*}(\Sigma)$ on its associated fan $\Sigma$ proven by Fulton and Strumfels in [FS97]. In [AHK18], Adiprasito, Huh and Katz bring these concepts together and prove that $A^{*}(M)$ and $MW_{*}(B(M))$ are dual to each other, in analogy to Poincaré duality for the toric case and the realizable case.
@@
where $\mathcal{L}$ is the lattice of non-empty flats of the matroid associated to $\mathcal{A}$ and $x_{F}$ is the element of the Chow ring of $\mathcal{W}$ associated to the flat $F$.
@@
indexed over all size $k$ multisubset chains =$\mathcal{F}=\{F_{r_{1}}\subseteq\dots\subseteq F_{r_{k}}\}$ of proper non-empty flats of $M$, with coefficients $C_{\mathcal{F}}$ on each monomial
@@
the staircase class of a matroid, which generalizes the formula provided in Section 3 to the case where $M$ is an arbitrary matroid. Interestingly, the staircase class satisfies a contraction-deletion formula as mentioned in Theorem 4.1, similar to that satisfied by Chern-Schwartz-MacPherson cycles. Since the proof of Theorem 4.1 is rather involved, the proof is contained in Section 5 instead of immediately after the statement of the theorem. The name “staircase” comes from the visual structure of the class when written out explicitly; in the case where $M=U_{3,3}$ with underlying set $E=\{0,1,2\}$,
@@
The proof of this statement uses the contraction-deletion formula, the interpretation of the Chow ring in terms of piecewise polynomials [Bri96] and induction on the number of elements of $M$, which completes the proof of Conjecture 5.2.4, Corollary 7.0.1 and Proposition 7.0.4 in [Man22].
Other work: In [Rau23], Rau provides a formula for the CSM cycle of a matroid $M$ in terms of the tropical polynomials cutting out the Bergman fan $1_{M}$, which is similar to Lemma 3.2 if we interpret the tropical polynomials as the sections whose zero loci are the components of the pullback of the hyperplane arrangement in the realizable case. In [BEST23], Berget, Eur, Spink and Tseng construct the framework of tautological classes of matroids, which generalizes the class of the universal subbundle and quotient bundle of the Grassmannian to an arbitrary matroid. Through this framework they obtain a formula for the inclusion of the group of Minkowski weights of a matroid $M$ into the Chow ring of the corresponding permutahedral variety and show that the Poincaré dual of $\operatorname{csm}(M)$ in the Chow ring of the permutahedral variety is $c(\mathcal{S}_{M})c_{|E|-\operatorname{rk}(M)}(\mathcal{Q}_{M})$. Although our proof of Theorem 4.2 does not use this framework, our result translates to the fact that for classes in $c_{|E|-\operatorname{rk}(M)}(\mathcal{Q}_{M})A^{*}(X_{E})$, $c(\mathcal{S}_{M})$ behaves as if its Chern roots are $\left\{-\sum_{\begin{subarray}{c}F\in\mathcal{L}\\ \operatorname{rk}(F)\geq r\end{subarray}}z_{F}|1\leq r\leq\operatorname{rk}(M)\right\}$ in $A^{*}(X_{E})$, via appendix III in [BEST23]. For specific details, see Section 4.3.
@@ 2 Background
Matroids arising in this way from a hyperplane arrangement in $\mathbb{P}^{d}_{\mathbb{C}}$ (or a vector arrangement in $\mathbb{C}^{d+1}$) are called $\mathbb{C}$-realizable. Note that there exist matroids which are not realizable.
In the realizable case, if $M=(\{0,\dots,k\},\operatorname{rk})$ is represented by the set of vectors $\{v_{0},\dots,v_{k}\}$ in $\mathbb{C}^{n}$, then $M\setminus i$ is represented by the set of vectors $\{v_{0},\dots,v_{k}\}\setminus\{v_{i}\}$ in $\mathbb{C}^{n}$ and $M/i$ is represented by $\{v_{0}+\braket{v_{i}},\dots,v_{k}+\braket{v_{i}}\}\setminus\{\braket{v_{i}}\}$ in $\mathbb{C}^{n}/\braket{v_{i}}$.
@@
Let $\pi:\mathcal{W}\rightarrow\mathbb{P}^{d}$ denote this blow-up map. $\pi$ is an isomorphism on the complement of the arrangement $C(\mathcal{A})$. There is a correspondence between the proper non-empty flats of the associated matroid to $\mathcal{A}$ and special divisors of $\mathcal{W}$, specifically, those that are (proper transforms of) exceptional divisors of the blowups described in the definition of $\mathcal{W}$. All collections of these divisors intersect transversely if their flats are comparable, and are disjoint otherwise. See [DCP96] for details.
@@
for all $i\in E$. This is analogous to the presentation of the Chow ring of the permutahedral variety mentioned in Appendix III.1 of [BEST23].
When a matroid is $\mathbb{C}$-realizable, arising from a hyperplane arrangement $\mathcal{A}$, its Chow ring is isomorphic to the Chow ring of the wonderful model $A^{*}(\mathcal{W})$ in the geometric sense, with each $x_{F}$ corresponding to the class of the divisor associated to the proper flat $F$, and $x_{E}=-\pi^{*}c_{1}(\mathcal{O}_{\mathbb{P}^{d}}(1))$, minus the pullback of the hyperplane class in $\mathbb{P}^{d}$.
@@
Let $M$ be the matroid in Example 2.1. Its Chow ring is presented by the quotient ring
@@
Let $E$ be a finite set and for $S\subset E$ define $e_{S}$ be the element of $\mathbb{R}^{E}$ given by $e_{S}(i)=1$ for all $i\in S$. Given a matroid $M=(E,\operatorname{rk})$, we define $\hat{B}(M)$, the affine Bergman fan of $M$ as the fan in $\mathbb{R}^{E}$ whose cones are of the form
@@
where =$\mathcal{F}=\{\emptyset\subsetneq F_{1}\subsetneq\dots\subsetneq F_{k}\subsetneq E\}$ is a flag of non-empty flats of $M$, containing the maximal flat $E$. We define the image of this fan under the projection $\mathbb{R}^{E}\to\mathbb{R}^{E}/\mathbb{R}e_{E}$ to be $B(M)$, the Bergman fan of $M$. Note that this projection maps $\hat{\sigma}_{\mathcal{F}\cup\{E\}}$ to the cone $\sigma_{\mathcal{F}}={\left\{\sum_{i=1}^{k}a_{i}e_{F_{i}}+\mathbb{R}e_{E}\mid a_{i}\geq 0\right\}}$, where $\mathcal{F}$ is a flag of proper non-empty flats.
@@
where $v_{\sigma/\tau}$ is the generator of $\sigma+\braket{\tau}$ in $N_{\mathbb{R}}/\braket{\tau}$. Given a fan $\Sigma$ of pure dimension $n$, we define $MW_{K}(\Sigma)$, the group $k$-dimensional Minkowski weights on $\Sigma$ to be the set of balanced weighted subfans of $\Sigma$ with addition given by addition on the weight functions. We define $MW_{*}(\Sigma)$ as the direct sum of the groups of $k$-dimensional Minkowski weights.
@@
Given a pure $n$-dimensional, balanced, unimodular fan $\Sigma$ in $N_{\mathbb{R}}$, $A^{k}(\Sigma)\cong MW_{n-k}(\Sigma)$. We denote the isomorphism by _$\_\cap 1_{\Sigma}$, where $1_{\Sigma}$ is the function $\sigma\mapsto 1$ for all $\sigma\in\Sigma(n)$. In the case where $\Sigma=B(M)$ for some matroid $M$, we simply write $1_{M}$ for $1_{B(M)}$. In particular, Remark 2.2 and the statements above imply that for all matroids $M$, $MW_{*}(\hat{B}(M))\cong MW_{*}(B(M))$.
@@ 3 A “Staircase” CSM Class Formula for Realizable Matroids
López de Medrano, Rincón, and Shaw defined an analogous CSM cycle for matroids, and proved that when a matroid is realizable over $\mathbb{C}$, its CSM cycle corresponds to the geometric CSM class $c_{SM}(\mathcal{W}\setminus\mathcal{A})$ of the complement of the arrangement in the wonderful model under the identification of $A_{*}(M)\cong A_{*}(\mathcal{W})$ (Theorem 3.1 in [LdMRS19]).
@@
In this section, we prove such a formula for the CSM class of the complement of the arrangement by geometric means, thus resolving the conjecture for all $\mathbb{C}$-realizable matroids.
@@
(Lemma 1.3 in [Alu09]) Let $X\subseteq Y$ be nonsingular varieties, and $\pi:\tilde{Y}\rightarrow Y$ the blow-up of Y along $X$, and $\tilde{X}$ the exceptional divisor. If $X$ is a complete intersection of hypersurfaces $Z_{1},\dots,Z_{d}$ meeting transversally in $Y$, then the Chern class of the tangent bundle of the blow up $\tilde{Y}$ is related to $c(T_{Y})$ by the formula
@@
The wonderful model $\mathcal{W}$ is constructed by a sequence of blow-ups of $\mathbb{P}^{d}$ at the intersection of hyperplanes determined by each flat $F\in\hat{\mathcal{L}}$. If we blow-up according to order by rank, highest to lowest (i.e. by dimension, lowest to highest), the center of each blow-up is the intersection of $\operatorname{rk}(F)$ many (pullbacks of) hyperplanes determined by $F$, minus the exceptional divisors already blown up in the intersection (corresponding to the flats $F^{\prime}$ which contain $F$). That is, when all blow-ups are complete, each of the $\operatorname{rk}(F)$ hyperplanes in locus of blow-up pulls back to
$\pi^{*}h-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}$ $=-x_{E}-\sum_{\begin{subarray}{c}F^{\prime}\in\hat{\mathcal{L}}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}$
$=-\sum_{\begin{subarray}{c}F^{\prime}\in\mathcal{L}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}$
in $A^{*}(\mathcal{W})$, where $h$ is the class of the hyperplane in $\mathbb{P}^{d}$, $c_{1}(\mathcal{O}_{\mathbb{P}^{d}}(1))$. By applying the blow-up formula (Lemma 3.1) according to each flat $F\in\hat{\mathcal{L}}$,
$c(T_{\mathcal{W}})=\prod_{F\in\hat{\mathcal{L}}}\left[\frac{(1+x_{F})(1-\sum_{\begin{subarray}{c}F^{\prime}\in\mathcal{L}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}-x_{F})^{\operatorname{rk}(F)}}{(1-\sum_{\begin{subarray}{c}F^{\prime}\in\mathcal{L}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}})^{\operatorname{rk}(F)}}\right]\pi^{*}c(T_{\mathbb{P}^{d}}).$
@@
$=\prod_{F\in\hat{\mathcal{L}}}\left[\frac{(1+x_{F})(1-\sum_{\begin{subarray}{c}F^{\prime}\in\mathcal{L}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}-x_{F})^{\operatorname{rk}(F)}}{(1+x_{F})(1-\sum_{\begin{subarray}{c}F^{\prime}\in\mathcal{L}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}})^{\operatorname{rk}(F)}}\right]\pi^{*}c(T_{\mathbb{P}^{d}})\cap[\mathcal{W}].$
By cancellation, and $\pi^{*}c(T_{\mathbb{P}^{d}})=(1+\pi^{*}h)^{d+1}=(1+\pi^{*}h)^{\operatorname{rk}(E)}=(1-x_{E})^{\operatorname{rk}(E)}$, we get
$c_{SM}(\mathcal{W}\setminus\mathcal{A})=\dots$ $=\prod_{F\in\hat{\mathcal{L}}}\left[\frac{(1-\sum_{\begin{subarray}{c}F^{\prime}\in\mathcal{L}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}}-x_{F})^{\operatorname{rk}(F)}}{(1-\sum_{\begin{subarray}{c}F^{\prime}\in\mathcal{L}\\ F^{\prime}\supsetneq F\end{subarray}}x_{F^{\prime}})^{\operatorname{rk}(F)}}\right](1-x_{E})^{\operatorname{rk}(E)}\cap[\mathcal{W}]$
@@
$\mathcal{F}$ must be a chain, since otherwise $\prod_{F\in\mathcal{F}}x_{F}=0$, since it corresponds to an empty intersection in $\mathcal{W}$. Choose a maximal flag $\{\emptyset=F_{0}\subsetneq F_{1}\subsetneq\dots\subsetneq F_{d}\subsetneq F_{d+1}=E\}$ (where $\operatorname{rk}(F_{i})=i$) containing the chain $\mathcal{F}$ as a multisubset. We may rewrite the monomial as
@@
For all $F$ not included in the flag, we may set $x_{F}=0$. This has no effect on $\prod_{F\in\mathcal{F}}x_{F}$, and only kills some other monomials. If $F$ is not in the flag, the fraction $\left(\frac{1-\sum_{F^{\prime}\supseteq F}x_{F^{\prime}}}{1-\sum_{F^{\prime}\supsetneq F}x_{F^{\prime}}}\right)^{\operatorname{rk}(F)}$ becomes $1$. Otherwise, when $F=F_{r}$ in the flag, it becomes $\left(\frac{1-\sum_{i=r}^{\operatorname{rk}(E)}x_{F_{i}}}{1-\sum_{i=r+1}^{\operatorname{rk}(E)}x_{F_{i}}}\right)^{r}$. The expression then simplifies by cancellation (we indicate the result of killing the non-flag divisors by $\mapsto$):
$\prod_{F\in\mathcal{L}}\left(\frac{1-\sum_{F^{\prime}\supseteq F}x_{F^{\prime}}}{1-\sum_{F^{\prime}\supsetneq F}x_{F^{\prime}}}\right)^{\operatorname{rk}(F)}$ $\mapsto\prod_{r=1}^{\operatorname{rk}(E)}\left(\frac{1-\sum_{i=r}^{\operatorname{rk}(E)}x_{F_{i}}}{1-\sum_{i=r+1}^{\operatorname{rk}(E)}x_{F_{i}}}\right)^{r}$
$=\prod_{r=1}^{\operatorname{rk}(E)}\left(\frac{1-(x_{F_{r}}+x_{F_{r+1}}+\dots+x_{E})}{1-(x_{F_{r+1}}+\dots+x_{E})}\right)^{r}$
$=\prod_{r=1}^{\operatorname{rk}(E)}\left(1-\sum_{i=r}^{\operatorname{rk}(E)}x_{F_{i}}\right)$
$=\prod_{r=1}^{\operatorname{rk}(E)}(1-(x_{F_{r}}+\dots+x_{E})).$
The coefficient of the monomial $x_{F_{r_{1}}}\cdots x_{F_{r_{k}}}$ may be deduced by counting how many times it the appears in the expansion111alternatively, we can use the Taylor’s Theorem to compute the coefficient by taking the appropriate partial derivatives at zero.. The count amounts to choosing from which factor each $x_{F_{r_{i}}}$ comes, then dividing to account for identical permutations. Note that each $-x_{F_{r_{i}}}$ appears as a term only in the first $r_{i}$ factors. There are $r_{1}$ possible choices for $x_{F_{r_{1}}}$, then $r_{2}-1$ many remaining choices for $x_{F_{r_{2}}}$, and so on (i.e. $r_{i}-i+1$ choices for $x_{F_{r_{i}}}$ once the previous $i-1$ variables have been chosen). Therefore, the coefficient on $x_{F_{r_{1}}}\cdots x_{F_{r_{k}}}$ is
@@
The formula conjectured by Fife and Rincón holds for $\mathbb{C}$-realizable matroids.